# DeepAAT: Deep Automated Aerial Triangulation for Fast UAV-based Mapping

### 02 Feb 2024 v1

@@ II-C Permutation Equivariant Layer
Theorem 1. [38] Take tensor $\mathbf{W}$ as input, the PEL with five unique parameters $h_{1}^{(d,o)},h_{2}^{(d,o)},...,h_{4}^{(d,o)}$ and $h_{5}^{(o)}$ could guarantee the output tensor $\mathbf{W}^{\prime}$ with size of $M\times N\times O$ exchangeable following a specific fully connected layer calculation rule in Eq. (4), where $d$ and $o$ are the indexes for the input and output feature channels, respectively.
@@
Inspired by the initial work proposed by Moran et al. [28], we also utilize PEL to extract exchangeable high-level geometry correlations from the feature track matrix $\mathbf{W}$. But different from Moran et al. [28], our proposed method takes into account not only the geometric features but also the spectral features of the feature tracks. Furthermore, the outliers in the feature track matrix are also automatically rejected to enhance the robustness of the results.
III System Overview
The proposed efficient UAV-based mapping system illustrated in Fig.2 is briefly introduced in this section. As most UAV controllers, such as PixHawk [39], depend on GPS for trajectory planning and tracking in survey applications, it is assumed that each UAV image is geotagged with GPS information provided by the UAV controller. Despite the Single Point Positioning (SPP) error of GPS potentially reaching 10 meters on the UAV, it can still serve as a useful guide for the image matching process, focusing on matching nearby images only, as demonstrated by [8]. To be compatible with distributed parallel processing and limit the GPU memory usage on one computing unit for large-scale UAV-based mapping, the proposed system exploits the hierarchical SfM scheme [17, 18], contains three components, namely, (1) image clustering, (2) DeepAAT, and (3) cluster merging.
@@ III System Overview
(1) Image clustering divides the entire image set into multiple subsets considering the 2D feature correspondences between images. By treating the complete image set as a scene graph $\mathbf{G(V,E)}$ [40], each image represents a vertex in $\mathbf{V}$, and an edge between two image vertices exists in $\mathbf{E}$ if the two images share feature correspondences. Setting the number of correspondences between images as the edge weight, $\mathbf{G(V,E)}$ is segmented using normalized cut [41] iteratively until the number of images in each subset is within a desired number $N_{subset}$. $N_{subset}$ can be set according to the GPU memory on each computing unit.
@@
Remark 1. The quantity of 2D feature correspondences between pairs of images typically serves as a crucial metric for evaluating the reliability of relative matching. In essence, the greater the number of 2D feature correspondences between image pairs, the higher their matching reliability, and conversely, the fewer the matches, the lower the reliability. Our goal is to achieve a strong level of mutual matching within each subset. Consequently, the objective of the normalized cut operation on $\mathbf{G(V,E)}$ is to minimize the sum of edge weights within the cut, while also ensuring a balanced distribution of elements in each subset to enhance the efficiency of computing for the following DeepAAT.
@@
(2) DeepAAT efficiently and robustly recovers camera poses and structural information within each cluster. The network structure and implementation details of DeepAAT will be described in Section IV.
@@
(3) Cluster merging conducts a global bundle adjustment of the entire images taking the camera poses from each subset as initial values, hence recovering the whole camera poses and structures in the targeting survey area. More specifically, after rejecting the outlier tracks identified by the DeepAAT, the remaining feature tracks are then re-triangulated using the initial camera poses resulting from DeepAAT. Then the global bundle adjustment is performed once to get the final result. Readers can refer to [11] for additional details on re-triangulation and global bundle adjustment.
IV Network Architecture of DeepAAT
@@
The network architecture of DeepAAT mainly consists of three parts, spatial-spectral feature aggregation module (Section IV-A), global consistency-based outlier rejecting module (Section IV-B), and pose decode module (Section IV-C), which are illustrated in Fig.3. The input of DeepAAT includes feature projection matrix $\mathbf{W}_{prj}$ constructed by reordering Eq.(3) with the shape of $M\times N\times 2$, SIFT feature descriptor matrix $\mathbf{W}_{des}$ with the shape of $M\times N\times 128$, and the geotag matrix $\mathbf{W}_{gps}$ with the shape of $M\times 3$. It’s important to highlight that the GPS-derived latitude, longitude, and altitude values are initially transformed into the East-North-Up (ENU) local coordinate system [42]. Following this transformation, they are further normalized to enhance network generativity. By feeding the input into the spatial-spectral feature aggregation module, a high-level embedded feature $\mathbf{W}_{em}$ is extracted for the downstream tasks. Then, the global consistency-based outlier rejecting module detects the outliers in the projection matrix $\mathbf{W}_{prj}$ and gives the confidence for each 2D observation using $\mathbf{W}_{outlier}$. Meanwhile, the pose decode module recovers the camera poses. The loss functions will be detailed in Section IV-D.
@@
IV-A Spatial-Spectral Feature Aggregation Module
Position Encoding: The coordinates within the feature projection matrix $\mathbf{W}_{prj}$ and the GPS matrix $\mathbf{W}_{gps}$ are both essential for the network to comprehend the spatial distribution. Position encoding [43] has been employed for both of these location-related pieces of information to enhance the network’s ability to distinguish relative positional relationships among input data, which is written as follows:
@@ IV-A Spatial-Spectral Feature Aggregation Module
where $x$ is the coordinate value in each dimension, $L$ is the coding level. Position encoding solely influences the last dimension of $\mathbf{W}_{prj}$ and $\mathbf{W}_{gps}$. As for the GPS matrix $\mathbf{W}_{gps}$, to ensure its dimensional consistency with $\mathbf{W}_{prj}$ and $\mathbf{W}_{des}$, the proposed method conducts a dimension expansion operation in the first dimension (image indexes), transforming it from a two-dimensional matrix into a three-dimensional one. Position encoding does not have learnable parameters, but it enhances the network’s ability to distinguish location information.
@@
Residual Permutation Equivariant Layer: The residual PEL employed in this paper consists of a consecutive pair of PEL, Instance Normalization (IN) [44], and Rectified Linear Unit (ReLU) combinations. Within the residual PEL, input and output data are summated through skip connections [45], serving two purposes: (1) ensuring stable gradient propagation within the network; and (2) facilitating the fusion of shallow layers, which contain more actual positional information, with deeper layers that offer more distinctive and discriminative feature information. It’s worth noting that IN and ReLU do not alter the permutation equivariant properties of PEL.
IV-B Global Consistency-based Outlier Rejecting Module
@@
Even with pair-wise Epipolar geometry verification, $\mathbf{W}_{prj}$ still contain a substantial number of outlier matches, which can significantly impact obtaining correct global triangulation results and ensuring the proper convergence of the BA. Therefore, the proposed method utilizes a global consistency outlier rejecting module, which, by leveraging global information from the embedded feature $\mathbf{W}_{em}$, ensures that the subsequent BA operates successfully.
@@
The global consistency-based outlier rejecting module primarily comprises three consecutive projection layers. Following the first two projection layers are the Context Normalization (CN) and ReLU activation functions. The CN primarily serves to integrate data, allowing the previous layer’s output to acquire corresponding context information in both camera poses and feature tracks. This aids the network in identifying outliers. Following the last projection layer is the Sigmoid. After passing through the Sigmoid function, the network’s output is a score matrix with values ranging from 0 to 1, with dimensions $M\times N$. For each non-empty point, the score represents the probability that each 2D feature is an inlier. The closer the score is to 1, the higher the confidence that it is an inlier. In the actual outlier detection process, with a given threshold $\tau_{outlier}$, the predicted scores greater than $\tau_{outlier}$ are considered as inliers, while those with scores less than $\tau_{outlier}$ are considered outliers.
IV-C Pose Decode Module
@@
The pose decode module utilizes the global information encoded by the spatial-spectral feature aggregation module to decode the camera’s position and orientation. The decoder first performs mean pooling on the input feature along the column dimension (feature tracks), which maps the embedded feature $\mathbf{W}_{em}$ with the shape of $M\times N\times O_{em}$ to $M\times O_{em}$. The reason for choosing mean pooling is that each camera observes a different number of 3D points in the input scene. By averaging the features observed by each camera, it can fairly represent the general characteristics of the scene observed by the $M$ cameras, regardless of the number of 3D points observed. This enables the decoder to obtain the global context information for each camera within the scene.
@@
IV-D Loss Function
The loss function for DeepAAT comprises three components: $\mathcal{L}_{outlier}$, $\mathcal{L}_{position}$, and $\mathcal{L}_{rotation}$, which written by:
@@ IV-D Loss Function
V Experiments
V-A Dataset
The experimental data, as depicted in Fig.4, was collected from an urban area including complex road networks, hills, construction sites, and buildings. A total of 4,992 UAV images were employed, subdivided into eight uniformed blocks labeled A to H. These images underwent preprocessing through SfM with Ground Control Points (GCPs) to establish reference data. Throughout the experiments, data from blocks A to G were used for training, while data from block H was employed as testing data.
@@ V-A Dataset
V-B Implementation Details
V-B1 Training sample generation
@@
V-B2 Data augmentation
To enhance the network’s generalizability, data augmentation is applied to the training data, focusing on two main aspects. Firstly, Gaussian noise with a mean of zero and a standard deviation of 0.01 was added to the input $\mathbf{W}_{prj}$. Secondly, given the limited amount of training data, and to augment and leverage the network’s permutation equivariance capability, random row and column permutations were applied to the training samples during training.
V-B3 Parameter settings
@@ V-B4 Evaluation criteria
where, $n_{2d}$ represents the number of 2D pixels in the scene, $m$ is the number of cameras, $n$ is the number of 3D points, and $x^{k}_{ij}$ represents the $k$th dimension of the coordinate of the $j$th 3D point observed by the $i$th image, $P_{i}^{k}$ represents the $k$th row of the $i$th camera matrix, $X_{j}$ represents the coordinate of the $j$th 3D point, $\hat{t}_{i}$ represents the reference value of camera position, $\widetilde{t}_{i}$ represents the predicted camera position, $\hat{R}_{i}$ represents the reference value of camera rotation, $\widetilde{R}_{i}$ represents the predicted camera rotation value, and $tr()$ represents the trace of the matrix (i.e. the sum of the main diagonal elements of the matrix).
@@ V-B5 Computation resources
V-C Results of scene segmenting and merging
@@
Figure 5: Image clustering result, where the number after ”_” represents the number of cameras included in the subset.
TABLE I: Outlier rejection result
@@
Table I shows that the average performance of the outlier rejection in the scene surpasses 0.95 across all four metrics, with a notable recall rate of 0.973. This indicates that about 97.3% of the pixels identified by the network as positive are indeed true positives. Such high accuracy is advantageous for the subsequent steps of global triangulation and BA. This also demonstrates that the global consistency-based outlier rejection module, as designed in this paper, is highly effective and applicable throughout the entire algorithmic process. Moreover, for clustered scenes, the network’s prediction time is under one second, highlighting the proposed network’s high operational efficiency in AAT tasks. Here, we predict all 8 scenarios through a single model loading, and except for the first scenario, the remaining 7 scenarios do not require reloading the model, resulting in a nearly tenfold reduction in time consumption.
TABLE II: Pose prediction result
@@
Table II demonstrates that across all eight clustered scenes, the predicted camera position error is consistently lower than the initial position error. The results show greater precision in predicting rotation, with an average error of less than 2°, which is highly beneficial for accurate subsequent adjustments. After BA, the average reprojection error across all scenes is less than 0.5 pixels, and the rotation error is under 0.1°. The visualized results, as depicted in Fig.6, further affirm the effectiveness of the algorithm.
@@
In the following, the Cluster Merging algorithm described in Section III is used to fuse the above 8 segmented scenes, and the results are shown in Fig.7 and Tab.III. From Fig.7, it can be seen that after using GPS information to globally align all segmented scenes, there is a significant offset between different scenes. After Cluster Merging, the inconsistency between scenes was effectively eliminated, resulting in globally consistent fusion results that were similar in appearance to the reference one. From Tab.III, it can be seen that the reprojection error of the scene has slightly decreased compared to the average reprojection error of the segmented scene, but in terms of position error and rotation error, it has slightly increased compared to the average result of the segmented scene. In addition, the reconstructed scene points have a slight increase compared to the 3D points of the reference scene. These results indicate that the proposed hierarchical AAT scheme can effectively complete large-scale AAT tasks.
@@
TABLE III: Cluster merging result

### 07 Apr 2024 v2

@@ 2.3 Permutation Equivariant Layer
Theorem 1. (Hartford et al., 2018) Take tensor $\mathbf{W}$ as input, the PEL with five unique parameters $h_{1}^{(d,o)},h_{2}^{(d,o)},...,h_{4}^{(d,o)}$ and $h_{5}^{(o)}$ could guarantee the output tensor $\mathbf{W}^{\prime}$ with size of $M\times N\times O$ exchangeable following a specific fully connected layer calculation rule in Eq. (4), where $d$ and $o$ are the indexes for the input and output feature channels, respectively.
@@
Inspired by the initial work proposed by Moran et al. (2021), we also utilize PEL to extract exchangeable high-level geometry correlations from the feature track matrix $\mathbf{W}$. But different from Moran et al. (2021), our proposed method takes into account not only the geometric features but also the spectral features of the feature tracks. Furthermore, the outliers in the feature track matrix are also automatically rejected to enhance the robustness of the results.
3 System Overview
The proposed efficient UAV-based mapping system illustrated in Fig.2 is briefly introduced in this section. As most UAV controllers, such as PixHawk (Meier et al., 2012), depend on GPS for trajectory planning and tracking in survey applications, it is assumed that each UAV image is geotagged with GPS information provided by the UAV controller. Despite the Single Point Positioning (SPP) error of GPS potentially reaching 10 meters on the UAV, it can still serve as a useful guide for the image matching process, focusing on matching nearby images only, as demonstrated by Schonberger and Frahm (2016). To be compatible with distributed parallel processing and limit the GPU memory usage on one computing unit for large-scale UAV-based mapping, the proposed system exploits the hierarchical SfM scheme (Chen et al., 2020; Xu et al., 2021), contains three components, namely, (1) image clustering, (2) DeepAAT, and (3) cluster merging.
@@ 3 System Overview
(1) Image clustering divides the entire image set into multiple subsets considering the 2D feature correspondences between images. By treating the complete image set as a scene graph $\mathbf{G(V,E)}$ (Zhu et al., 2018), each image represents a vertex in $\mathbf{V}$, and an edge between two image vertices exists in $\mathbf{E}$ if the two images share feature correspondences. Setting the number of correspondences between images as the edge weight, $\mathbf{G(V,E)}$ is segmented using normalized cut (Shi and Malik, 2000) iteratively until the number of images in each subset is within a desired number $N_{subset}$. $N_{subset}$ can be set according to the GPU memory on each computing unit.
@@
Remark 1. The number of 2D feature correspondences between pairs of images typically serves as a crucial metric for evaluating the reliability of relative matching. In essence, the greater the number of 2D feature correspondences between image pairs, the higher their matching reliability, and conversely, the fewer the matches, the lower the reliability. Our goal is to achieve a strong level of mutual matching within each subset. Consequently, the objective of the normalized cut operation on $\mathbf{G(V,E)}$ is to minimize the sum of edge weights within the cut, while also ensuring a balanced distribution of elements in each subset to enhance the computational efficiency for the following DeepAAT.
@@
(2) DeepAAT efficiently and robustly recovers camera poses and structural information within each cluster. The network structure and implementation details of DeepAAT will be described in Section 4.
@@
(3) Cluster merging conducts a global bundle adjustment of the entire images taking the camera poses from each subset as initial values, hence recovering the whole camera poses and structures in the targeting survey area. More specifically, after rejecting the outlier tracks identified by the DeepAAT, the remaining feature tracks are then re-triangulated using the initial camera poses resulting from DeepAAT. Then the global bundle adjustment is performed once to get the final result. Readers can refer to Triggs et al. (2000) for additional details on re-triangulation and global bundle adjustment.
4 Network Architecture of DeepAAT
@@
The network architecture of DeepAAT mainly consists of three parts, spatial-spectral feature aggregation module (Section 4.1), global consistency-based outlier rejecting module (Section 4.2), and pose decode module (Section 4.3), which are illustrated in Fig.3. The input of DeepAAT includes feature measurement matrix $\mathbf{W}_{mes}$ constructed by reordering Eq.(3) with the shape of $M\times N\times 2$, SIFT feature descriptor matrix $\mathbf{W}_{des}$ with the shape of $M\times N\times 128$, and the geotag matrix $\mathbf{W}_{gps}$ with the shape of $M\times 3$. It’s important to note that the GPS-derived latitude, longitude, and altitude values are initially transformed into the East-North-Up (ENU) local coordinate system (Shin and El-Sheimy, 2002). Following this transformation, they are further normalized to enhance network generativity. By feeding the input into the spatial-spectral feature aggregation module, a high-level embedded feature $\mathbf{W}_{em}$ is extracted for the downstream tasks. Then, the global consistency-based outlier rejecting module detects the outliers in the measurement matrix $\mathbf{W}_{mes}$ and gives the confidence for each 2D observation using $\mathbf{W}_{outlier}$. Meanwhile, the pose decode module recovers the camera poses. The loss functions will be detailed in Section 4.4.
@@
4.1 Spatial-Spectral Feature Aggregation Module
Position Encoding: The coordinates within the feature measurement matrix $\mathbf{W}_{mes}$ and the GPS matrix $\mathbf{W}_{gps}$ are both essential for the network to comprehend the spatial distribution. Position encoding (Mildenhall et al., 2021) has been employed for both of these location-related pieces of information to enhance the network’s ability to distinguish relative positional relationships among input data, which is written as follows:
@@ 4.1 Spatial-Spectral Feature Aggregation Module
where $x$ is the coordinate value in each dimension, $L$ is the coding level. Position encoding solely influences the last dimension of $\mathbf{W}_{mes}$ and $\mathbf{W}_{gps}$. As for the GPS matrix $\mathbf{W}_{gps}$, to ensure its dimensional consistency with $\mathbf{W}_{mes}$ and $\mathbf{W}_{des}$, the proposed method conducts a dimension expansion operation in the first dimension (image indexes), transforming it from a two-dimensional matrix of M × 3 into a three-dimensional one of M × N × 3. Position encoding does not have learnable parameters, but it enhances the network’s ability to distinguish location information.
@@
Residual Permutation Equivariant Layer: The residual PEL employed in this paper consists of a consecutive pair of PEL, Instance Normalization (IN) (Ulyanov et al., 2016), and Rectified Linear Unit (ReLU) combinations. Within the residual PEL, input and output data are summated through skip connections (He et al., 2016), serving two purposes: (1) ensuring stable gradient propagation within the network; and (2) facilitating the fusion of shallow layers, which contain more actual positional information, with deeper layers that offer more distinctive and discriminative feature information. It’s worth noting that IN and ReLU do not alter the permutation equivariant properties of PEL.
4.2 Global Consistency-based Outlier Rejecting Module
@@
Even with pair-wise Epipolar geometry verification, $\mathbf{W}_{mes}$ still contain a substantial number of outlier matches, which can significantly impact obtaining correct global triangulation results and ensuring the proper convergence of the BA. Therefore, the proposed method utilizes a global consistency outlier rejecting module, which, by leveraging global information from the embedded feature $\mathbf{W}_{em}$, ensures that the subsequent BA operates successfully.
@@
The global consistency-based outlier rejecting module primarily comprises three consecutive projection layers, which are fully connected (FC) layers that change the number of channels for non-empty data in sparse matrices, i.e. $Proj:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{\prime}}$. Following the first two projection layers are the Context Normalization (CN) and ReLU activation functions. The CN primarily serves to integrate data, allowing the previous layer’s output to acquire corresponding context information in both camera poses and feature tracks. This aids the network in identifying outliers. Following the last projection layer is the Sigmoid. After passing through the Sigmoid function, the network’s output is a score matrix with values ranging from 0 to 1, with dimensions $M\times N$. For each non-empty point, the score represents the probability that each 2D feature is an inlier. The closer the score is to 1, the higher the confidence that it is an inlier. In the outlier detection process, with a given threshold $\tau_{outlier}$, the predicted scores greater than $\tau_{outlier}$ are considered as inliers, while scores less than $\tau_{outlier}$ are considered outliers.
4.3 Pose Decode Module
@@
The pose decode module utilizes the global information encoded by the spatial-spectral feature aggregation module to decode the camera’s position and orientation. The decoder first performs mean pooling on the input feature along the column dimension (feature tracks), which maps the embedded feature $\mathbf{W}_{em}$ with the shape of $M\times N\times O_{em}$ to $M\times O_{em}$. The reason for choosing mean pooling is that each camera observes a different number of 3D points in the input scene. By averaging the features observed by each camera, the network can fairly represent the general characteristics of the scene observed by the $M$ cameras, regardless of the number of 3D points observed. This enables the decoder to obtain the global context information for each camera within the scene.
@@
4.4 Loss Function
The loss function for DeepAAT comprises three components: $\mathcal{L}_{outlier}$, $\mathcal{L}_{position}$, and $\mathcal{L}_{rotation}$, which are written by:
@@ 4.4 Loss Function
5 Experiments
5.1 Dataset
The experimental data, as depicted in Fig.4, was collected from an urban area including complex road networks, hills, construction sites, and buildings. A total of 4,992 UAV images were employed, subdivided into eight uniformed blocks labeled A to H. These images underwent preprocessing through SfM with high-precision Ground Control Points (GCPs) to establish reference data. Throughout the experiments, data from blocks A to G were used for training, while data from block H was employed as testing data. During dataset preparation, feature points that can be successfully matched but do not appear in the final AAT result are labeled as outliers.
@@ 5.1 Dataset
5.2 Implementation Details
5.2.1 Training sample generation
@@
5.2.2 Data augmentation
To enhance the network’s generalizability, data augmentation is applied to the training data, focusing on two main aspects. Firstly, Gaussian noise with a mean of zero and a standard deviation of 0.01 was added to the input $\mathbf{W}_{mes}$. Secondly, given the limited amount of training data, and to augment and leverage the network’s permutation equivariance capability, random row and column permutations were applied to the training samples in advance.
5.2.3 Sparse matrix
Because the number of scene points that can be observed in each image is limited, there will be a large number of zero elements in the measurement matrix $\mathbf{W}_{mes}$ and the descriptor matrix $\mathbf{W}_{des}$. Therefore, these matrices are implemented in the form of sparse matrices in the code to improve the processing efficiency of the network.
5.2.4 Parameter settings
@@ 5.2.5 Evaluation criteria
where $n_{2d}$ represents the number of 2D pixels in the scene, $m$ is the number of cameras, $n$ is the number of 3D points, and $x^{k}_{ij}$ represents the $k$th dimension of the coordinate of the $j$th 3D point observed by the $i$th image, $P_{i}^{k}$ denotes the $k$th row of the $i$th camera matrix, $X_{j}$ denotes the coordinate of the $j$th 3D point, $\hat{t}_{i}$ is the reference value of camera position, $\widetilde{t}_{i}$ is the predicted camera position, $\hat{R}_{i}$ is the reference value of camera rotation, $\widetilde{R}_{i}$ is the predicted camera rotation value, and $tr()$ denotes the trace of the matrix (i.e. the sum of the main diagonal elements of the matrix).
@@ 5.2.6 Computation resources
5.3 Results of scene segmenting and merging
@@
Figure 5: Image clustering result, where the number after "_" represents the number of cameras included in the subset.
Table 1: Outlier rejection result
@@
Table 1 shows that the average performance of the outlier rejection in the scene surpasses 0.95 across all four metrics, with a notable recall rate of 0.973. This indicates that about 97.3% of the pixels identified by the network as positive are indeed true positives. Such high accuracy is advantageous for the subsequent steps of global triangulation and BA. This also demonstrates that the global consistency-based outlier rejection module, as designed in this paper, is highly effective and applicable throughout the entire algorithmic process. Moreover, for clustered scenes, the network’s prediction time is under one second, highlighting the proposed network’s high operational efficiency in AAT tasks. Here, we predict all 8 scenarios through a single model loading, and except for the first scenario, the remaining 7 scenarios do not require reloading the model, resulting in a nearly tenfold reduction in time consumption.
Table 2: Pose prediction result
@@
Table 2 demonstrates that across all eight clustered scenes, the predicted camera position error is consistently lower than the initial position error. The results show greater precision in predicting rotation, with an average error of less than 2°, which is highly beneficial for accurate subsequent adjustments. After BA, the average reprojection error across all scenes is less than 0.5 pixels, and the rotation error is under 0.1°. The visualized results, as depicted in Fig.6, further affirm the effectiveness of the algorithm.
@@
In the following, the Cluster Merging algorithm described in Section 3 is used to fuse the above 8 segmented scenes, and the results are shown in Fig.7 and Tab.3. From Fig.7, it can be seen that after using GPS information to globally align all segmented scenes, there is a significant offset between different scenes. After Cluster Merging, the inconsistency between scenes was effectively eliminated, resulting in globally consistent fusion results that were similar in appearance to the reference one. From Tab.3, it can be seen that the reprojection error of the scene has slightly decreased compared to the average reprojection error of the segmented scene, but in terms of position error and rotation error, it has slightly increased compared to the average result of the segmented scene. In addition, the reconstructed scene points have a slight increase compared to the 3D points of the reference scene. These results indicate that the proposed hierarchical AAT scheme can effectively complete large-scale AAT tasks.
@@
Table 3: Cluster merging result